OpenStudy (anonymous):

the number of telephones in a certain town is 48,000. if this represents 12.8 telephones per 100 of population, the population of this town to the nearest thousand is what? A. 128,000 B. 375,000 C.378,000 D. 566,000 E. 560,000

OpenStudy (anonymous):

So you could do proportions\[\frac{ 12.8 }{ 100 }=\frac{ telephones }{ population }\]So you would have\[\frac{ 12.8 }{ 100 }=\frac{ 48,000 }{ population(p) }\] So cross multiply and solve for p or population\[12.8p=4,800,000\]then divide by 12.8 to get population\[p=\frac{ 4,800,000 }{ 12.3 }=375,000\]So your answer would be choice B.

OpenStudy (anonymous):

Hope this helps and Good Luck :)

OpenStudy (anonymous):

oh wait my bad you don't divide by 12.3 you divide by 12.8

OpenStudy (anonymous):

Typo you still get 375,000 so still choice B

OpenStudy (anonymous):

thank you